OpenStudy (anonymous):

5/6 + 1/3 / 6/7 simplest form

OpenStudy (anonymous):

@EmmaMink

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

5/6+1/3= 1 1/6 1 1/6 ? 6/7= 1 13/36

OpenStudy (anonymous):

I meant to type 1 1/6 / 6/7 = 1 13/36. sorry it was a typo